OpenStudy (anonymous):

please help!! 1 medal and fan Identify the property and equation that satisfies the following statement: The solution of an equation is x = –2. Addition Property of Equality; x + 6 = 8 Substitution Property of Equality; x + 8 = 6 Division Property of Equality; x + 8 = 6 Subtraction Property of Equality; x + 6 = 8

OpenStudy (anonymous):

Substitution Property of Equality; x + 8 = 6

OpenStudy (anonymous):

identity is subtraction property of equality equation is x + 8 = 6

OpenStudy (anonymous):

take the equation x + 8 = 6. subtract 8 from both sides using the subtraction property of equality x + 8 - 8 = 6 - 8 simplify x + 0 = -2 x = -2

OpenStudy (anonymous):

THANK YOU ITS appreciated

OpenStudy (anonymous):

yw